OpenStudy (anonymous):

Felix exclaims that his quadratic with a discriminant of -1 has no real solutions. Felix then puts down his pencil and refuses to do any more work. Create an equation with a negative discriminant. Then explain to Felix, in calm and complete sentences, how to find the solutions, even though they are not real. (10 points)

OpenStudy (anonymous):

\(x^2+1=0\)

OpenStudy (anonymous):

This has a negative discriminant.. \(a=1\), \(b=0\), \(c=1\) \(\sqrt{b^2-4(a)(c)}=\sqrt{0^2-4(1)(1)}=\sqrt{-4}\)

OpenStudy (anonymous):

But you can solve this easily

OpenStudy (anonymous):

\(x^2 = -1\) \(x = \pm \sqrt{-1}\) \(x=\pm i\), since \(i = \sqrt{-1}\)